# Subsystem Information Capacity in Random Circuits and Hamiltonian Dynamics

### v2

@@ Abstract
In this study, we explore the information capacity of open quantum systems, focusing on the effective channels formed by the subsystem of random quantum circuits and quantum Hamiltonian evolution. By analyzing the subsystem information capacity, which is closely linked to quantum coherent information of these effective quantum channels, we uncover a diverse range of dynamical and steady behaviors depending on the types of evolution. Therefore, the subsystem information capacity serves as a valuable tool for studying the intrinsic nature of various dynamical phases, such as integrable, localized, thermalized, and topological systems. We also reveal the impact of different initial information encoding schemes on information dynamics including one-to-one, one-to-many, and many-to-many. To support our findings, we provide representative examples for numerical simulations, including random quantum circuits with or without mid-circuit measurements, random Clifford Floquet circuits, free and interacting Aubry-André models, and Su-Schrieffer-Heeger models. Those numerical results are further quantitatively explained using the effective statistical model mapping and the quasiparticle picture in the cases of random circuits and non-interacting Hamiltonian dynamics, respectively.
I Introduction
Quantum information can present rich and intriguing dynamical behaviors in different non-equilibrium phases. The principles that govern the propagation of information often defy intuition, contrasting with the more familiar laws of matter and energy transfer. Despite their counterintuitive nature, these principles are of vital importance for advancing our comprehension of quantum thermalization Deutsch (1991); Srednicki (1994); Rigol et al. (2008), quantum chaos Maldacena et al. (2016); Xu et al. (2020); Dowling et al. (2023), quantum gravity Hayden and Preskill (2007); Brown et al. (2016), quantum computation and error correction Lloyd (2000); Choi et al. (2020); Fan et al. (2021); Li and Fisher (2021).
Entanglement Amico et al. (2008) and its growth Calabrese and Cardy (2005); Kim and Huse (2013); Liu and Suh (2014); Alba and Calabrese (2017); Ho and Abanin (2017); Mezei and Stanford (2017); Nahum et al. (2017); Bertini et al. (2022) constitute fundamental building blocks for modern condensed matter physics and usually offer valuable insights from information perspectives for the analysis of quantum many-body phases. Out-of-time correlators (OTOC) Maldacena et al. (2016); Lewis-Swan et al. (2019); Xu and Swingle (2024), as another commonly used information probe, also attracts intensive academic interest owing to its multifaceted phenomena, experimental relevance and relations with information scrambling Hayden and Preskill (2007); Sekino and Susskind (2008); Lashkari et al. (2013); Hosur et al. (2016) and holographic duality.
@@ I Introduction
We demonstrate the equivalence between the subsystem information capacity investigated here and the concept of quantum coherent information Schumacher and Nielsen (1996) for the effective subsystem channels in some cases. This equivalence further connects the subsystem information capacity to the single-shot quantum channel capacity Lloyd (1997); DiVincenzo et al. (1998); Barnum et al. (1998); Shor (2003); Devetak and Winter (2003); Gyongyosi et al. (2018) in quantum communication. The monotonically decreasing nature of quantum coherent information with the application of any quantum channels indicates that only the effective channels that maintain the initial value of quantum coherent information can admit perfect quantum error correction Colmenarez et al. (2023). We also discuss the similarities and differences between subsystem information capacity and Holevo information, serving as bounds for quantum and classical information transmission capacities, respectively. The relation between OTOC and SIC is also explored.
In this work, we examine the information aspects of effective subsystem channels in a variety of systems that hold significant experimental relevance or offer clear analytical insights within certain limits. Our investigation encompasses quantum circuits with brickwall two-qubit gates, incorporating randomness in spatial or temporal dimensions, with or without mid-circuit measurements. Additionally, we study information dynamics in time evolution controlled by both non-interacting and interacting Aubry-Andrée (AA) models and Su-Schrieffer-Heeger (SSH) models.
This paper is organized as follows. In Sec. II we introduce the setups in this work including the models, the encoding schemes, and the information metrics of interest. In Sec. III, we present the results from random quantum circuits and apply the entanglement membrane picture to quantitatively understand these results. In Sec. IV, we present numerical results for the time evolution of several one-dimensional representative fermionic Hamiltonians and understand the results with the quasiparticle picture in non-interacting limits. In Sec. V, we discuss possible future directions and conclude. Several technical details and supplemental results are presented in the Appendixes.
II General setups
In this section, we introduce the setups utilized in this study in three parts: models that describe the dynamics including quantum circuits and time-independent Hamiltonian evolution, schemes for information encoding at the initial stage, and the probes employed to characterize information dynamics. An overview of our general setups and encoding schemes is provided in Fig. 1.
@@ II General setups
II.1 Models
@@
In our analysis of circuit models, we focus on one-dimensional random quantum circuits Fisher et al. (2023), featuring a structured arrangement of even-odd brickwall layers with random two-qudit gates drawn from Haar or Clifford ensembles. We also consider the cases with mid-circuit measurements where the properties of corresponding dynamics are averaged over different measurement history trajectories Li et al. (2018, 2019); Skinner et al. (2019a); Chan et al. (2019). Such mid-circuit measurements can significantly influence the global information structure and dynamics, offering a contrast to the local unitary dynamics.
We also consider the random Clifford circuit setup where the layers of two-qudit gates are time-periodic, leading to what we term random Clifford Floquet dynamics. Previous studies reported that 1D random Clifford Floquet systems show Anderson localization behaviors Farshi et al. (2022, 2023), while 1D random Haar Floquet circuits tend to thermalize Sünderhauf et al. (2018). Numerically, we consider systems with qubits $d=2$, while in the theoretical analysis, we consider systems composed of $L$ $d$-qudits with both large $d$ limit and finite $d$ corrections ensuring a comprehensive understanding of the dynamics.
In terms of Hamiltonian dynamics, we consider one-dimensional spinless fermionic Hamiltonians. One representative example we investigated is the Aubry-André model without or with many-body interactions Harper (1955); Aubry and André (1980); Iyer et al. (2013); Khemani et al. (2017); Zhang and Yao (2018, 2019). This model is well-known for its Anderson localization (many-body localization) phase and extended (thermal) phase in non-interacting (interacting) cases. The Hamiltonian of the AA model reads as:
$H=$ $J\sum_{j=1}^{L}(c_{j+1}^{\dagger}c_{j}+h.c.)+U\sum_{j=1}^{L}n_{j}n_{j+1}$
$+$ $\sum_{j=1}^{L}V_{j}c_{j}^{\dagger}c_{j},$ (1)
We also investigate the interplay between information and topological properties Qi and Zhang (2011), where SSH model is utilized Su et al. (1980):
$H=-\sum_{j=1}^{L}(1-(-1)^{j}\Delta)c_{j}^{\dagger}c_{j+1}+h.c.$ (2)
The quench dynamics under investigation is described by the Hamiltonian evolution unitary $e^{-iHt}$ with $H$ defined in Eq. (1) for AA models or Eq. (2) for SSH models.
II.2 Encoding schemes
@@
In the meantime, other system qudits $\bar{E}$ remain in the same product state as described above for initial states. The one-to-one encoding scheme is different from the typical information scrambling settings in defining tripartite mutual information (TMI) or connecting OTOC with Rényi entropy, where other system qubits $\bar{E}$ are also initialized with fully mixed states Hosur et al. (2016). Therefore, our encoding schemes are different from conventional information scrambling settings and more qubit-efficient for real quantum hardware experiments (see Sec. II.3.3 for detailed discussions).
@@ II.3 Probes
II.3.1 Subsystem information capacity
In our settings, as discussed in Sec. II.2, the system qubits are divided into $E$ and $\bar{E}$ for the initial states. We further categorize the system qubits into intervals $A$ and $\bar{A}$ for the output states at time $t$. Generally, the number of freedoms of the region $E$ and the region $A$ are not necessarily the same. We identify the effective channel $\mathcal{E}_{sub}$ formed by the subsystem evolution mapping $E$ to $A$ as
@@ II.3.1 Subsystem information capacity
where $U$ describes the dynamics given by the quantum circuit or Hamiltonian evolution, $\rho_{E}=\text{Tr}_{R}(\rho_{ER})$ for initial pure state $\rho_{ER}$ under given encoding scheme and $\rho_{\bar{E}}$ is the initial state for subsystem $\bar{E}$. Such an effective channel is formed by the subsystem of some dynamics after tracing out the complementary system as given in Eq. (8), we call it the “subsystem channel" that can be regarded as an evolution with some form of dissipation induced by the complementary subsystem. Such subsystem channels are interesting and explicitly useful for exploring the intrinsic properties of a system as they are formed by tracing out homogeneous physical freedoms instead of prototypical models where the freedoms for the system and environment bath are heterogeneous.
@@
There are a variety of indicators for evaluating information propagation and protection, such as quantum coherent information, OTOC, tripartite mutual information, and Holevo information. In the following discussions, we clarify the differences and similarities between SIC and these quantities.
II.3.2 Relation with quantum coherent information
@@
For an abitrary quantum channel $\mathcal{E}$ that maps the state from $\rho$ defined on system $Q$ to $\rho^{\prime}=\mathcal{E}(\rho)$ defined on system $Q^{\prime}$ (the number of freedom can be different for $\rho$ and $\rho^{\prime}$, i.e. $|Q|\neq|Q^{\prime}|$), we introduce the ancilla system $R$ and the pure state $|\Psi_{RQ}\rangle$ such that it reproduce the input state $\rho$ on $Q$ as $\rho=\text{Tr}_{R}|\Psi_{RQ}\rangle\langle\Psi_{RQ}|$. The quantum coherent information is defined in terms of the output state $\rho^{\prime}_{RQ^{\prime}}=(\mathcal{I}\otimes\mathcal{E})(|\Psi_{RQ}\rangle\langle\Psi_{RQ}|)$:
@@
The advantage of quantum coherent information lies in that it has a data processing inequality Schumacher and Nielsen (1996), which states that the quantity cannot increase with the application of quantum channels. The initial coherent quantum information is the entropy of the input state $S(\rho_{E})=S(\rho_{A})$ which sets the upper limit of the coherent quantum information of any output state. As quantum error correction operations are also quantum channels, the monotonicity of this quantity suggests that perfect information protection and decoding are only feasible with channels that preserve the initial value of coherent information ($I(A:R)=2$ for unitary evolution $U$). Consequently, this measure serves as a faithful reflection of the channel’s information capacity. Indeed, this quantity is used to lower bound the single-shot quantum channel capacity Lloyd (1997); DiVincenzo et al. (1998); Barnum et al. (1998); Shor (2003); Devetak and Winter (2003); Gyongyosi et al. (2018), which gives the maximal amount of quantum information that can be reliably transferred via the quantum channel.
II.3.3 Relation with OTOC
OTOC is an important quantity that can diagnose information scrambling. It is defined as $F=\text{Tr}(W_{i}(t)V_{j}W_{i}(t)V_{j})$ for infinite temperature system. If the two local operators $W$ and $V$ average over all Pauli operators on two regions $\bar{A}$ and $E$, the corresponding averaged OTOC is related to the Rényi-2 entropy $S^{(2)}_{AR}$ Hosur et al. (2016). However, $S^{(2)}_{AR}$ here is defined in a different setting from our case, where the input state in $\bar{E}$ is a full mixed state $\rho_{\bar{E}}\propto I$. Specifically, all system qubits are entangled with reference qubits individually, leaving the input to the unitary $U$ a fully mixed state on $E\cup\bar{E}$. On the contrary, we employ the input state on $\bar{E}$ as a pure state $\rho_{\bar{E}}=|\psi_{\bar{E}}\rangle\langle\psi_{\bar{E}}|$. In sum, our settings differ from conventional settings for mutual information and are highly qubit-efficient for numerical simulations and experiments.

### v3

@@ Abstract
In this study, we explore the information capacity of open quantum systems, focusing on the effective channels formed by the subsystem of random quantum circuits and quantum Hamiltonian evolution. By analyzing the subsystem information capacity, which is closely linked to quantum coherent information of these effective quantum channels, we uncover a diverse range of dynamical and steady behaviors depending on the types of evolution. Therefore, the subsystem information capacity serves as a valuable tool for studying the intrinsic nature of various dynamical phases, such as integrable, localized, thermalized, and topological systems. We also reveal the impact of different initial information encoding schemes on information dynamics including one-to-one, one-to-many, and many-to-many. To support our findings, we provide representative examples for numerical simulations, including random quantum circuits with or without mid-circuit measurements, random Clifford Floquet circuits, free and interacting Aubry-André models, and Su-Schrieffer-Heeger models. These numerical results are further quantitatively explained using the effective statistical model mapping and the quasiparticle picture in the cases of random circuits and non-interacting Hamiltonian dynamics, respectively.
1 Introduction
Quantum information can present rich and intriguing dynamical behaviors in different non-equilibrium phases. The principles that govern the propagation of information often defy intuition, contrasting with the more familiar laws of matter and energy transfer. Despite their counterintuitive nature, these principles are of vital importance for advancing our comprehension of quantum thermalization [1, 2, 3], quantum chaos [4, 5, 6], quantum gravity [7, 8], quantum computation and error correction [9, 10, 11, 12].
Entanglement [13] and its growth [14, 15, 16, 17, 18, 19, 20, 21] constitute fundamental building blocks for modern condensed matter physics and usually offer valuable insights from information perspectives for the analysis of quantum many-body phases. Out-of-time correlators (OTOC) [4, 22, 23], as another commonly used information probe, also attract intensive academic interest owing to its multifaceted phenomena, experimental relevance and relations with information scrambling [7, 24, 25, 26] and holographic duality.
@@ 1 Introduction
We demonstrate the equivalence between the subsystem information capacity investigated here and the concept of quantum coherent information [27] for the effective subsystem channels in some cases. This equivalence further connects the subsystem information capacity to the single-shot quantum channel capacity [28, 29, 30, 31, 32, 33] in quantum communication. The monotonically decreasing nature of quantum coherent information with the application of any quantum channels indicates that only the effective channels that maintain the initial value of quantum coherent information can admit perfect quantum error correction [34]. We also discuss the similarities and differences between subsystem information capacity and Holevo information, serving as bounds for quantum and classical information transmission capacities, respectively. The relation between OTOC and SIC is also explored.
In this work, we examine the information aspects of effective subsystem channels in a variety of systems that hold significant experimental relevance or offer clear analytical insights within certain limits. Our investigation encompasses quantum circuits with brickwall two-qubit gates, incorporating randomness in spatial or temporal dimensions, with or without mid-circuit measurements. Additionally, we study information dynamics in time evolution controlled by both non-interacting and interacting Aubry-André (AA) models and Su-Schrieffer-Heeger (SSH) models.
This paper is organized as follows. In Sec. 2 we introduce the setups in this work including the models, the encoding schemes, and the information metrics of interest. In Sec. 3, we present the results from random quantum circuits and apply the entanglement membrane picture to quantitatively understand these results. In Sec. 4, we present numerical results for the time evolution of several one-dimensional representative fermionic Hamiltonians and understand the results with the quasiparticle picture in non-interacting limits. In Sec. 5, we discuss possible future directions and conclude. Several technical details and supplemental results are presented in the Appendices.
2 General setups
In this section, we introduce the setups utilized for the definition of SIC in terms of models that describe the dynamics including quantum circuits and time-independent Hamiltonian evolution, schemes for information encoding at the initial state, and the probes employed to characterize information dynamics. An overview of our general setups and encoding schemes is provided in Fig. 1. We also discuss the relations between SIC and other information metrics and the unique advantages for SIC.
@@ 2 General setups
SIC provides a direct measure of how effectively a specific spatial subsystem $A$ can preserve or recover quantum information initially encoded in the region $E$ after undergoing time evolution $U$. Its core strengths include:
  • Direct Information-Theoretic Meaning: SIC is fundamentally linked to the quantum coherent information [27, 28], quantifying the capacity of the effective subsystem channel to transmit quantum information. This provides a clear operational interpretation relevant to quantum communication, memory, and error correction.
  • Rich Spatio-Temporal Information: SIC, defined as a function of subsystem size $x=|A|$ and time $t$, provides a detailed map $MI(x,t)$ revealing not only the speed of information propagation but also its spatial distribution and structure, particularly in the long-time limit.
  • Broad Applicability and Sensitivity: SIC effectively distinguishes diverse dynamical phases, including thermalizing, integrable, Anderson localized, many-body localized (MBL), and topological phases, revealing characteristic “fingerprints” in its late-time spatial profile (see Fig. 2).
  • Experimental Advantages: Compared to standard protocols for OTOCs or Tripartite Mutual Information (TMI), measuring SIC often requires simpler state preparation (fewer entangled qubits) and measurement protocols (only one forward time evolution $U$), rendering it potentially more feasible on near-term quantum hardware.
2.1 Models
@@
In our analysis of circuit models, we focus on one-dimensional random quantum circuits [35], featuring a structured arrangement of even-odd brickwall layers with random two-qudit gates drawn from Haar or Clifford ensembles. We also consider the cases with mid-circuit measurements where the properties of corresponding dynamics are averaged over different measurement history trajectories [36, 37, 38, 39]. Such mid-circuit measurements can significantly influence the global information structure and dynamics, offering a contrast to the local unitary dynamics.
We also consider the random Clifford circuit setup where the layers of two-qudit gates are time-periodic, leading to what we term random Clifford Floquet dynamics. Previous studies reported that 1D random Clifford Floquet systems show Anderson localization behaviors [40, 41], while 1D random Haar Floquet circuits tend to thermalize [42].
In terms of Hamiltonian dynamics, we consider one-dimensional spinless fermionic Hamiltonians. These Hamiltonians $H$ can exhibit thermal, localization, topology features and the details of these Hamiltonians are reported in Sec. 4. Consequently, the quench dynamics under investigation is described by the Hamiltonian evolution unitary $e^{-iHt}$.
2.2 Encoding schemes
@@
In the meantime, other system qudits $\bar{E}$ remain in the same product state as described above for initial states. The one-to-one encoding scheme is different from the typical information scrambling settings in defining tripartite mutual information (TMI) or connecting OTOC with Rényi entropy, where other system qubits $\bar{E}$ are also initialized with fully mixed states [26]. Therefore, our encoding schemes are different from conventional information scrambling settings and more qubit-efficient for real quantum hardware experiments (see Appendix D for detailed discussions).
@@ 2.3 SIC Probes
In our settings, as discussed in Sec. 2.2, the system qubits are divided into $E$ and $\bar{E}$ for the initial states. We further categorize the system qubits into intervals $A$ and $\bar{A}$ for the output states at time $t$. Generally, the number of freedoms of the region $E$ and the region $A$ are not necessarily the same. We identify the effective channel $\mathcal{E}_{sub}$ formed by the subsystem evolution mapping $E$ to $A$ as
@@
where $U$ describes the dynamics given by the quantum circuit or Hamiltonian evolution, $\rho_{E}=\text{Tr}_{R}(\rho_{ER})$ for initial pure state $\rho_{ER}$ under given encoding scheme and $\rho_{\bar{E}}$ is the initial state for subsystem $\bar{E}$. Such an effective channel is formed by the subsystem of some dynamics after tracing out the complementary system as given in Eq. (6), we call it the “subsystem channel” that can be regarded as an evolution with some form of dissipation induced by the complementary subsystem. Such subsystem channels are interesting and explicitly useful for exploring the intrinsic properties of a system as they are formed by tracing out homogeneous physical freedoms instead of prototypical models where the freedoms for the system and environment bath are heterogeneous.
@@
2.4 Late-time SIC as “fingerprints” for dynamical phases
Following the presentation of our general frameworks, we now highlight one of its most powerful applications: using the late-time spatial profile of SIC as a characteristic “fingerprint” to identify and distinguish different quantum dynamical phases. While the temporal evolution $MI(x,t)$ reveals the dynamics of information propagation, the focus here is on the steady-state or long-time limit, $MI(x,t\to\infty)$. This quantity, when plotted as a function of the subsystem size fraction $x/L=L_{A}/L$ (typically for the one-to-one encoding scheme with $E$ initially at the center or boundary), reflects the stable spatial distribution achieved by the initially localized quantum information at long time limits.
Our central finding, concisely illustrated in Fig. 2, is that this late-time SIC curve exhibits qualitatively distinct and characteristic shapes depending fundamentally on the nature of the underlying dynamics (e.g., chaotic, integrable, localized, topological). This allows SIC to serve as an insightful diagnostic tool. By treating the evolution $U$ as a “black box”, measuring the $MI(x,t\to\infty)$ profile provides a remarkably clear means to discern intrinsic properties of the system’s dynamical phase. The shape of this curve encapsulates how effectively information initially encoded in $E$ can be recovered from subsystems $A$ of varying size in the long run, revealing crucial details about information transport and memory within the system and is potentially helpful in quantum error correction context.
This “fingerprint” capability offers potentially richer information than simpler metrics. For instance, while steady-state entanglement entropy often follows volume or area laws for broad classes of phases, it might not fully capture the structural differences within those classes. As summarized in Table 1, different phases that might all exhibit volume-law entanglement (e.g., thermalizing chaotic systems vs. certain interacting integrable systems) can display markedly different late-time SIC profiles. SIC probes the structure and accessibility of specific initially encoded quantum information distributed across spatial partitions.
The origin of these distinct fingerprints lies in the fundamental mechanisms governing information transport in different phases.
  • In chaotic/thermalizing systems, information scrambles rapidly throughout the system, leading to a near-complete loss of recoverable information from any small local subsystem $|A|<L/2$, resulting in a profile (Fig. 2(d)) until $A$ encompasses a significant fraction of the system.
  • Anderson localized systems completely halt information transport. Information initially in $E$ remains localized, yielding an SIC profile as in Fig. 2(b).
  • Many-body localized systems exhibit an intermediate behavior with slow (logarithmic) information spreading but incomplete delocalization. This results in a unique SIC profile (Fig. 2(f)) distinct from both thermal and Anderson phases, reflecting partial information kept across larger distances.
  • Systems with localized modes, such as topological edge states, can trap information near the corresponding modes. If $E$ is placed near such a mode, the SIC profile can exhibit plateaus reflecting this trapping (Fig. 2(c)).
  • Integrable systems can show varied behaviors depending on the nature of their quasiparticle excitations (e.g., compare Fig. 2(a) for free fermions vs. (d) for certain interacting cases).
Therefore, the late-time SIC profile $MI(x,t\to\infty)$ provides a nuanced and powerful lens through which to classify and understand the fundamental nature of quantum dynamics, complementing and extending insights gained from traditional entanglement and scrambling measures. It paves the way for a deeper comprehension of how information behaves in complex quantum systems, which is essential for both fundamental physics and applications in quantum information processing.
Figure 2: Late-time subsystem information capacity with one-to-one encoding scheme as a function of the subsystem ratio $L_{A}/L$, highlighting the qualitative differences in behavior across various dynamical phases. Each panel presents the x-axis as $L_{A}/L$ and the y-axis as the average late-time SIC, with the initial value 2. (a) Extended phase of non-interacting fermion, with some integrable models following similar trends. (b) Anderson localization phase of non-interacting fermion. Anderson localization phase induced by random Clifford Floquet circuit also follows similar curves. (c) Topological phase with edge states with the entangled qubit at the spatial boundary. (d) Chaotic phases including random circuits and thermal phases of non-integrable interacting systems. Notably, some non-thermal phases of integrable interacting systems also display comparable behavior. (e) Noise-resilient phase in monitored random circuits. (f) Many-body localization phases.
Phase Entanglement growth Steady-state entanglement Steady-state SIC
Chaotic phase (Sec. 3.2.1, 4.3.1) Linear growth Volume law (Haar value) (d) in Fig. 2
Many-body localization (Sec. 4.3.1) Logarithmic growth Volume law (smaller than Haar value) (f) in Fig. 2
Anderson localization (Sec. 3.2.2, 4.2) No growth Area law (b) in Fig. 2
Noise-resilient phase (Sec. 3.2.1) Linear growth Volume law (smaller than Haar value) (e) in Fig. 2
Free fermion extended phase (Sec. 4.2) Linear growth Volume law (a) in Fig. 2
Interacting integrable systems (Sec. 4.3.2) Linear growth Volume law (d) in Fig. 2
Topological edge states (Sec. 4.4) Linear growth Volume law (smaller than Haar value) (c) in Fig. 2
SWAP circuits (Sec. 4.2) No growth Area law (a) in Fig. 2
CNOT circuits (Sec. 3.2.3) No growth Area law (possible long-range entanglement) Fig. 11
2.5 Relation with other information metrics
@@
There are a variety of indicators for evaluating information propagation and protection, such as quantum coherent information, OTOC, tripartite mutual information, and Holevo information. In the following discussions, we clarify the differences and similarities between SIC and these quantities and compares the scope and advantages of SIC relative to those metrics.
2.5.1 Relation with quantum coherent information
@@
For an arbitrary quantum channel $\mathcal{E}$ that maps the state from $\rho$ defined on system $Q$ to $\rho^{\prime}=\mathcal{E}(\rho)$ defined on system $Q^{\prime}$ (the number of freedom can be different for $\rho$ and $\rho^{\prime}$, i.e. $|Q|\neq|Q^{\prime}|$), we introduce the ancilla system $R$ and the pure state $|\Psi_{RQ}\rangle$ such that it reproduces the input state $\rho$ on $Q$ as $\rho=\text{Tr}_{R}|\Psi_{RQ}\rangle\langle\Psi_{RQ}|$. The quantum coherent information is defined in terms of the output state $\rho^{\prime}_{RQ^{\prime}}=(\mathcal{I}\otimes\mathcal{E})(|\Psi_{RQ}\rangle\langle\Psi_{RQ}|)$:
@@
The advantage of quantum coherent information lies in that it has a data processing inequality [27], which states that the quantity cannot increase with the application of quantum channels. The initial coherent quantum information is the entropy of the input state $S(\rho_{E})=S(\rho_{A})$ which sets the upper limit of the coherent quantum information of any output state. As quantum error correction operations are also quantum channels, the monotonicity of this quantity suggests that perfect information protection and decoding are only feasible with channels that preserve the initial value of coherent information ($I(A:R)=2$ for unitary evolution $U$). Consequently, this measure serves as a faithful reflection of the channel’s information capacity. Indeed, this quantity is used to lower bound the single-shot quantum channel capacity [28, 29, 30, 31, 32, 33], which gives the maximal amount of quantum information that can be reliably transferred via the quantum channel.
2.5.2 Relation with OTOC
OTOC is an important quantity that can diagnose information scrambling. It is defined as $F=\text{Tr}(W_{i}(t)V_{j}W_{i}(t)V_{j})$ for infinite temperature systems. If the two local operators $W$ and $V$ average over all Pauli operators on two regions $\bar{A}$ and $E$, the corresponding averaged OTOC is related to the Rényi-2 entropy $S^{(2)}_{AR}$ [26]. However, $S^{(2)}_{AR}$ here is defined in a different setting from our case, where the input state in $\bar{E}$ is a full mixed state $\rho_{\bar{E}}\propto I$. Specifically, all system qubits are entangled with reference qubits individually, leaving the input to the unitary $U$ a fully mixed state on $E\cup\bar{E}$. On the contrary, we employ the input state on $\bar{E}$ as a pure state $\rho_{\bar{E}}=|\psi_{\bar{E}}\rangle\langle\psi_{\bar{E}}|$. In sum, our settings differ from conventional settings for mutual information and are highly qubit-efficient for numerical simulations and experiments.